OpenStudy (anonymous):

HELPPP!!! 1.rectangular based pyramid with height of 5 feet and a base with dimensions of 6 feet and 3 feet A.-27 B-30 C-33 D-28

OpenStudy (anonymous):

Volume = base* height What's the height?

OpenStudy (anonymous):

THE H = 5 FT.

OpenStudy (anonymous):

ok correct. Now the base is the area of the rectangle. The area of a rectangle is area=w*l w=6ft l=3ft

OpenStudy (anonymous):

Can you find the base?

OpenStudy (anonymous):

18 ??

OpenStudy (anonymous):

@dpaInc help!

OpenStudy (anonymous):

careful.... you're looking for the volume of a pyrmid, not a prism...

OpenStudy (anonymous):

volume of a pyrmid = (1/3)(area base)(height)

OpenStudy (anonymous):

ok so volume \[Volume=\frac{1}{3}*base*h\]

OpenStudy (anonymous):

@lashae1995 , 18 is the area of the base.

OpenStudy (anonymous):

1/3 * 18 * h ?

OpenStudy (anonymous):

height is given as 5